Suppose a population consists of the numbers 2,3,5,8 and 12. What is the mean of the




population? (Clue: Point Estimation)





1
Expert's answer
2022-02-21T09:37:59-0500

Population mean:

"\mu=\dfrac{\sum _ix_i}{n}=\dfrac{2+3+5+8+12}{5}=6"